Find the number of combinations and permutations of 4 letters taken from the word EXAMINATION

Grade:12

2 Answers

Askiitians Expert Ankit Jain- IIT Bombay
13 years ago

Hello Aditya,

Its a nice question you have posted.

The word examination consists of letters in following frequencies.

E=1; X=1; A=2; M=1; N=2; T=1; O=1; I=2;

For picking up the 4 letters and arrange them into words woul require us to consider 3 cases.

 

CASE I:  2 same letters each of 2 kinds

 THe possibilities are:

AANN

AAII

NNII

Each word can be arranged in 4!/2!*2! ways. So for three cases total ways = 3* 4!/2!*2! =18 ways.

 

CASE II: 2 same letters of one kind

 The possbilities are:

AA _ _

NN _ _

II _ _

Now consider the case with AA _ _. For selecting the next two letter you can select maximum of 1  N and 1 I. So you need to select 2 letters from 7 left. ( Total =11. 2 left for AA , 1 N and 1 I left). 

So say for AA  _ _ case total number of permutations = C(7,2) 4!/2!

So total permutations for all the three cases = 3 * C(7,2) 4!/2! = 378 ways

 

CASE III: No repetition of any letter

Total no of permutation for this can be directly written as = C(8,4) * 4! =1680 ways

 

So total number of ways = 1680 + 378 + 18 =2076 ways

 

Hope this solves your query.

Please feel free to post as many doubts on our disucssion forum as you can.  We are all IITians and here to help you in your IIT JEE preparation.

All the best!!


Regards

Askiitians Expert

Ankit Jain- IIT Bombay

Devraj konwar
6 years ago
how many combination and permutations can be made with the letters of the word PARABOCA, taken three at a time?